OpenStudy (anonymous):

A material decays at a rate of 1.1% per year. If you start with 250 grams of the material, in how many years will there be 100 grams remaining? Round to one decimal place. What would the equation look like?

OpenStudy (anonymous):

Is the answer 100=250(98.9)^t?

OpenStudy (anonymous):

dm/dt = -0.011 m m(t) = m(0) exp (-0.011 t) t in years get t from solving m(t)/m(0) = 100/250 = exp(-0.011 t) by taking logarithms

OpenStudy (anonymous):

Oh okay. So is the answer 100=250(98.9)^t?

OpenStudy (anonymous):

ln (100/250) = - 0.011 t t = [ln (100/250)]/[-0.011] = 83.3 years

OpenStudy (anonymous):

But is the equation correct? 100=250(89.9)^t?

OpenStudy (anonymous):

eys

OpenStudy (anonymous):

on no

OpenStudy (anonymous):

to decrease a number by\(1.1\%\) multiply it by \(100\%-1.1\%=98.9\%=.989\)

OpenStudy (anonymous):

so it should be \[100=250(.989)^t\] solve for \(t\)

OpenStudy (anonymous):

work with numbers, not percents

OpenStudy (anonymous):

Ohhh okay THANK YOU! Alright so now how do I solve that... I get confused once I am at .4=(.989)^t

OpenStudy (anonymous):

@satellite73